# Flaming-hot Initiation with Regular Execution Sampling for Large Language Models

### v1

@@ Abstract
Guanlin Liu, Renjie Zheng, Wenlei Shi, Chen Dun, Zheng Wu, Xing Jin, Lin Yan ByteDance {guanlin.liu, renjie.zheng, wenlei.shi}@bytedance.com {chen.dun, zheng.wu1, jinxing.9, neil}@bytedance.com
@@ 2 Related Works
Our work is closely related to CoT-decoding Wang and Zhou (2024), which uncovers the CoT-paths by enumerating over the top-k alternative tokens and aggregating the responses by scoring the decoded responses with confidence on the final answer. However, our approach differs in three key aspects: (1) we introduce a differentiable sampling method that can be directly integrated with existing inference and training frameworks, (2) we focus on improving model performance in scenarios with a sandbox checker, where aggregating responses is less data-efficient, and (3) our method operates without assumptions about the prompts, even when a chain of thought (CoT) is included, extending beyond the scope of CoT-decoding.

### v2

@@ Abstract
Guanlin Liu, Renjie Zheng, Wenlei Shi, Chen Dun, Zheng Wu, Xing Jin, Lin Yan ByteDance Inc. {guanlin.liu, renjie.zheng, wenlei.shi}@bytedance.com {chen.dun, zheng.wu1, jinxing.9, neil}@bytedance.com
@@ 2 Related Works
Our work is closely related to CoT-decoding Wang and Zhou (2024), which uncovers the CoT-paths by enumerating over the top-k alternative tokens and aggregating the responses by scoring the decoded responses with confidence on the final answer. However, our approach differs in three key aspects: (1) we introduce a differentiable sampling method that can be directly integrated with existing inference and training frameworks, (2) we focus on improving model performance in scenarios with a sandbox checker, where aggregating responses is less data-efficient, and (3) our method operates without assumptions about the prompts, even when a chain of thought (CoT) is included, extending beyond the scope of CoT-decoding. Prior work has explored dynamic temperature allocation during processing Chang et al. (2023); Zhu et al. (2023); Zhang et al. (2024). However, these approaches restrict the temperature values to the range of $[0,1]$, which contrasts sharply with the extreme temperature settings employed in our algorithm.
Orthogonal to our work, numerous efforts in the NLP domain have proposed prompt-based algorithms to enhance the diversity of generated responses Naik et al. (2023); Huang et al. (2023); Wu et al. (2023); Zhou et al. (2024). In contrast, our method focuses on scenarios where the prompt remains fixed, improving generation diversity by controlling the sampling process without relying on domain-specific modifications to the prompt. Notably, our approach is complementary to these works and can be combined with prompt-based methods to further enhance diversity, particularly in cases where generating a larger number of diverse samples for the same question is required.
@@ 3.2 Experiments
Dataset Model Sampling Pass@1 Pass@5 Pass@10 Pass@20
GSM8K Qwen Reg $66.95\pm 0.09$ $89.33\pm 0.06$ $92.58\pm 0.04$ $94.48\pm 0.03$
FIRE $65.65\pm 0.12$ $92.92\pm 0.04$ $95.58\pm 0.03$ $97.20\pm 0.02$
Deepseek Reg $87.02\pm 0.06$ $93.64\pm 0.03$ $95.06\pm 0.03$ $96.02\pm 0.02$
FIRE $85.85\pm 0.07$ $94.86\pm 0.03$ $96.76\pm 0.02$ $97.60\pm 0.02$
MATH Qwen Reg $35.94\pm 0.05$ $59.32\pm 0.03$ $66.04\pm 0.03$ $71.19\pm 0.02$
FIRE $39.64\pm 0.05$ $63.34\pm 0.03$ $69.75\pm 0.03$ $74.77\pm 0.02$
Deepseek Reg $51.03\pm 0.04$ $64.45\pm 0.03$ $68.58\pm 0.02$ $71.80\pm 0.01$
FIRE $48.79\pm 0.05$ $65.39\pm 0.02$ $70.25\pm 0.02$ $73.99\pm 0.02$
Table 6: Average and standard deviation of pass rate results for different models on different datasets.
@@
Is the improvement stable?
In previous sections and the appendix, we demonstrated that FIRE consistently enhances performance across various models, datasets, and hyperparameter configurations. Here, we further substantiate these findings by evaluating their statistical significance. Specifically, for each question, we leverage a pool of 40 pre-collected samples and randomly select different subsets of $n$ samples, repeating this process 10 times. The results, summarized in Table 6, reveal minimal standard deviations across the datasets, underscoring the robustness and statistical stability of the improvements.